OpenStudy (anonymous):

Simplify the expression. 3 log 3 (6x – 4)

OpenStudy (anonymous):

Simplify the expression. 3 log 3 (6x – 4)

OpenStudy (alfie):

\[\begin{array}{l} {\rm{3 log 3 }}\left( {{\rm{6x }}-{\rm{ 4}}} \right)\\ OR\\ {\rm{3 lo}}{{\rm{g}}_3}\left( {{\rm{6x }}-{\rm{ 4}}} \right) \end{array}\] In any case... \[a\log b = \log {b^a}\]